# Hisham Cholakkal1  Rao Muhammad Anwer1,2  Salman Khan1,3  Fahad Khan1,4

### v2

@@ 1 Introduction
Large multimodal models (LMMs) that jointly handle vision and language have made substantial progress in recent years. They excel at tasks like image captioning [Lu2025OmniCaptionerOCA, Jung2025VisualANA, Peng2025MattersTFA, Li2025DescribeEarthDAA], visual question answering [Yang2025MAGICVQAMAA, Jiang2025FastOSA, Zhou2025HierarchicalVRA, Peng2025MVCoReMVA], and multimodal reasoning [Yang2025R1OnevisionAGA, Zhang2025R1VLLTA, Wang2025VisualPRMAEA, Peng2025SkyworkRPA, thawakar2025llamav] by leveraging large-scale pre-training and fine-tuning on annotated datasets. However, despite these advances, two fundamental dependencies remain: (i) the dependence on human-curated annotations or metadata [Hudson2019GQAANA, Zhao2023MMICLEVA, AlTahan2024UniBenchVRA, Dai2024HumanVLMFFA, Liu2024MMDUAMA], and (ii) the dependency on externally defined reward models or evaluators (e.g., human-labeled correctness, knowledge-distilled reward networks) [wang2025vision, cai2025llavakdframeworkdistillingmultimodal, shu2024llavamod, 11094175, li2024self]. These dependencies pose scalability limitations, hinder domain generalization, and restrict deployment in settings where annotations or reward models are unavailable.
@@
Recent research in large language models (LLMs) explores the idea of self-evolving models that not only solve tasks but also generate the tasks they learn from. For instance, few works [chen2025self, chen2025multi] aim to improve base LLM by generating its own questions and attempting to answer them, using the agreement among multiple sampled answers as a self-reward signal. Such a technique can be further improved [chen2025multi] by splitting the model into interacting roles such as Proposer, Solver, and Judge that co-exist and train themselves through reinforcement learning. This research direction is pushed further [zhang2025darwin] by incorporating open-ended scheme where the system incorporates an open-ended evolution to continually refine internal behaviors. However, these research efforts are largely constrained to the language-only domain, where reasoning is explicit and correctness can be inferred through textual consistency or symbolic execution. Moreover, these approaches still either rely on implicit evaluators (e.g., external Judge module, execution feedback) or explicit evaluators, implying that the learning signal is not yet fully free of external priors.
@@
In the multimodal (LMM) domain, few recent works aim to improve LMMs through self-rewarding training signals, thereby reducing human supervision. For instance, the work of  [li2025self] proposes a self-rewarding reasoning strategy where the model decomposes each task into a perception step and a reasoning step, and then evaluates with internal bootstrapping. Similarly, ViPER [zhang2025viper] introduces a closed-loop framework that gradually strengthens visual perception ability through image-level and instance-level reconstruction, combined with reinforcement style updates driven by internally synthesized data. While these approaches represent progress in self-improving LMMs, they still rely on structured intermediate signals such as reconstruction objectives, heuristic quality filtering, or semantic similarity-based scoring. Consequently, their self-evolution remains partially guided by external priors rather than emerging purely from the model’s own reasoning dynamics.
@@ 2 Related Work
In the context of large language models (LLMs), several recent works [chen2025self, shafayat2025can, chen2025multi] aim to self-improve the model capabilities (e.g., scientific reasoning) without significant human annotations. The work of [chen2025self] proposes an approach to improve the implicit reasoning capabilities of the base model through self-play and majority-vote reward signals. Sheikh shafayat2025can explores self-consistency as a training reward within an reinforcement learning (RL) framework, demonstrating performance on synthetic reasoning task. Recently, the work of [chen2025multi] introduces a self-play based method where three roles (Proposer, Solver, and Judge) are instantiated from the same base LLM and trained via RL to construct questions, answers, and the evaluation inter-play between them. While achieving promising results, these approaches assume a certain form of verifiable environment in the form of a judge [chen2025multi] or heuristic and discrete rewards (e.g., majority vote) [chen2025self, shafayat2025can] which can likely struggle with model collapse.
In the context of large multimodal models (LMMs), li2025self introduces a self-rewarding technique to improve visual understanding by decomposing it into perception and language-reasoning via explicit perception reward by the LMM itself. However, the self-reward mechanism relies on the SFT cold-start dataset to learn the desired format. The work of zhang2025viper leverages image-level and instance-level reconstructions in a two-stage RL loop to iteratively improve perceptual capabilities. However, this image-and instance-level reconstruction of meta-data relies on an external model (e.g., OmniGen2 [wu2025omnigen2] and Qwen-Image [wu2025qwen]). Additionally, a recent work Vision-Zero [wang2025vision], trains LMMs through a multi-agent social deduction game (“Who is the Spy?”) played directly on images. The framework introduces adversarial roles (spy vs. civilians), iterative clue and voting rounds, and alternates between self-play and verifiable reinforcement learning (RLVR) depending on whether the identity of the spy is known. Although Vision-Zero does not require explicit human annotations, it relies on constructing image pairs with subtle, controlled differences produced through automated rendering, image editing, or regeneration pipelines relying on external generative systems (GPT-4o [gpt4o] and Gemini [comanici2025gemini]).
@@ 3.1 Unsupervised Self-Evolving LMM
In this work, we design an unsupervised self-evolving LMM approach (EvoLMM) motivated by the recent self-question-based technique (SQLM) [chen2025self] in the LLM domain. The SQLM improves an LLM by generating its own questions and reinforcing answer agreement. However, SQLM relies on a discrete majority-vote reward, which we empirically show to be sub-optimal in the multimodal setting where the LMM is required to ground its questions based on the input image and perform complex multimodal reasoning (e.g., visual math reasoning). We observe that early-stage Solver outputs over images are highly variable, leading to frequent zero-reward updates and unstable optimization. This prevents the model from making gradual progress when using discrete majority-vote reward.
@@
Training Optimization. In our framework, the Proposer and Solver are trained jointly in an online reinforcement-learning loop. Since no ground-truth labels or external scoring functions are available, both policies rely entirely on internally generated reward signals. To ensure stable learning, we adopt a REINFORCE policy gradient [williams1992simple] with two key components: (i) moving-average baselines to reduce variance, and (ii) token-level KL regularization to monitor deviation from the pretrained base LMM.
@@ 3.2 EvoLMM Training Insights
To present further insights into our continuous self-consistency reward-based training, we conduct an experiment to post-train Qwen2.5-VL-7B [qwen25vlreport] on ChartQA [chartqa] training images (without any question-answer pairs or metadata). As shown in Figure 5, our self-consistency reward yields a graded score proportional to the empirical agreement among Solver outputs. Even partial agreement (e.g., 2 out of 5 consistent samples) produces a meaningful positive reward. Furthermore, our continuous feedback reduces collapse and accelerates Solver improvement (see Figure 4). The continuous Solver reward changes smoothly and slightly dips at partial consensus, while the Proposer peaks at moderate difficulty (see step 5122 in Figure 5).
@@ 4 Experiments
Vision-Zero (CLEVR) [wang2025vision] $84.24$ $68.43$ $23.96$ $43.86$ $80.35$ $82.64$ $88.50$ $51.44$
Qwen2.5-VL-7B (Baseline) [qwen25vlreport] $84.00$ $68.46$ $23.91$ $43.78$ $80.44$ $82.61$ $88.30$ $51.11$
@@ 4.1 Dataset
Our training uses only raw images, with no question–answer pairs or metadata. We sample roughly 1,000 images each from six widely used benchmarks such as ChartQA [chartqa], AI2D [ai2d], InfographicVQA [infovqa], PlotQA [plotqa], ChartX [chartx], and Geometry3K [geovqa] totaling $\sim$6k images. These datasets cover charts, plots, scientific diagrams, and geometric figures, providing diverse visual contexts for the Proposer to generate grounded mathematical questions. This diversity encourages the Solver to develop stable, structured reasoning as the system self-evolves using internal consistency alone.
@@ 4.3 Evaluation
We evaluate our self-evolving framework on a broad suite of mathematical and visually grounded reasoning benchmarks, including ChartQA [chartqa], MathVista [mathvista], MathVision [mathvision], MathVerse [mathverse], InfoGraphic-VQA [infovqa], AI2D [ai2d], ScienceQA [scienceqa], and MMMU [mmmu]. These datasets test a range of capabilities from chart interpretation and symbolic math to diagram understanding—and all metrics follow their standard accuracy protocols. We use the official evaluation splits and apply identical inference settings across models, ensuring that performance differences arise solely from our self-evolving training rather than task-specific tuning. All evaluations are run on AMD MI250X GPUs using the lmms-eval framework [lmmseval], with HuggingFace Transformers v4.38 and bfloat16 precision for consistency with training.
@@
Qwen2.5-VL-7B (Base) [qwen25vlreport] $84.00$ $68.46$ $23.91$ $43.78$ $80.44$ $82.61$ $88.30$ $51.11$
@@
InternVL3-8B-Instruct (Base) [zhu2025internvl3] $82.40$ $65.2$ $25.36$ $31.62$ $68.77$ $83.19$ $97.77$ $52.78$
@@
Gemma3-12B-It (Base) [team2025gemma] $55.64$ $60.13$ $24.53$ $28.96$ $50.69$ $79.05$ $83.89$ $48.11$
@@
Llama-3.2-11B-Vision-Instruct (Base) [meta2024llama3_2] $29.24$ $46.59$ $23.47$ $37.23$ $56.69$ $46.44$ $56.87$ $47.93$
@@ 4.4 Main Results
In Table 1, we compare our proposed EvoLMM framework against two baselines: (1) the original pretrained Qwen2.5-VL-7B model [qwen25vlreport] without any additional training, and (2) a direct multimodal adaptation [chen2025self] that uses a discrete majority-vote reward during training. Qwen2.5-VL-7B (Baseline). The baseline model performs well on visually grounded reasoning tasks that require direct reading or surface-level interpretation (e.g., ChartQA trend identification, InfoGraphic-VQA object recognition). However, the model’s performance drops on tasks requiring multi-step inference or symbolic reasoning (e.g., MathVista, MathVerse), indicating that pretrained multimodal alignment alone is insufficient for deeper reasoning. We also report results from recent Vision-Zero [wang2025vision], whose self-play formulation bears partial similarity to ours through autonomous multi-agent learning. However, Vision-Zero leverages synthetic image pairs generated using external models such as GPT-4o [gpt4o] and Gemini [comanici2025gemini], introducing implicit supervision that falls outside our unsupervised setting. Vision-SR1 [li2025self] differs entirely from our constraints, as it relies on paired question–answer supervision across both perception and reasoning stages.
@@ 4.5 Comparison Across Training Settings
We evaluate the self-evolving framework under three fine-tuning strategies such as LoRA [lora], QLoRA [qlora], and full-parameter fine-tuning—while keeping the Proposer–Solver pipeline and continuous reward formulation identical across settings (see Table 2).
@@ 4.6 Self Evolution Across Different Backbones
To evaluate the generality of our approach, we apply the same Proposer–Solver training setup to several large multimodal models, including Qwen3-VL-8B-Instruct [qwen25vlreport], InternVL3-8B [zhu2025internvl3], Gemma-3-12B-it [team2025gemma], and Llama-3.2-11B-Vision-Instruct [meta2024llama3_2]. As shown in Table 3, in all cases, the models are trained without question–answer supervision, using only the proposed continuous self-consistency reward. We observe consistent improvements of roughly $2-3\%$ across all backbones, confirming that our method is architecture-agnostic and does not rely on specific dataset alignments or handcrafted supervision. Models with stronger visual grounding (e.g., InternVL3-8B [zhu2025internvl3]) show faster convergence in early training, while those with stronger linguistic reasoning (e.g., Gemma-3-12B [team2025gemma]) display larger late-stage gains as solver reasoning stabilizes. These results demonstrate that continuous self-reward effectively generalizes across diverse LMM families. This shows that our reward mechanism enhances structured visual reasoning rather than surface-level pattern matching, providing a scalable and backbone-independent path toward autonomous multimodal reasoning.

### v3

@@ 1 Introduction
Large multimodal models (LMMs) that jointly handle vision and language have made substantial progress in recent years. They excel at tasks like image captioning [22, 14, 29, 16], visual question answering [42, 13, 51, 28], and multimodal reasoning [43, 46, 37, 30, 34] by leveraging large-scale pre-training and fine-tuning on annotated datasets. However, despite these advances, two fundamental dependencies remain: (i) the dependence on human-curated annotations or metadata [11, 50, 1, 8, 19], and (ii) the dependency on externally defined reward models or evaluators (e.g., human-labeled correctness, knowledge-distilled reward networks) [36, 4, 32, 12, 17]. These dependencies pose scalability limitations, hinder domain generalization, and restrict deployment in settings where annotations or reward models are unavailable.
@@
Recent research in large language models (LLMs) explores the idea of self-evolving models that not only solve tasks but also generate the tasks they learn from. For instance, few works [5, 6] aim to improve base LLM by generating its own questions and attempting to answer them, using the agreement among multiple sampled answers as a self-reward signal. Such a technique can be further improved [6] by splitting the model into interacting roles such as Proposer, Solver, and Judge that co-exist and train themselves through reinforcement learning. This research direction is pushed further [45] by incorporating open-ended scheme where the system incorporates an open-ended evolution to continually refine internal behaviors. However, these research efforts are largely constrained to the language-only domain, where reasoning is explicit and correctness can be inferred through textual consistency or symbolic execution. Moreover, these approaches still either rely on implicit evaluators (e.g., external Judge module, execution feedback) or explicit evaluators, implying that the learning signal is not yet fully free of external priors.
@@
In the multimodal (LMM) domain, few recent works aim to improve LMMs through self-rewarding training signals, thereby reducing human supervision. For instance, the work of  [18] proposes a self-rewarding reasoning strategy where the model decomposes each task into a perception step and a reasoning step, and then evaluates with internal bootstrapping. Similarly, ViPER [47] introduces a closed-loop framework that gradually strengthens visual perception ability through image-level and instance-level reconstruction, combined with reinforcement style updates driven by internally synthesized data. While these approaches represent progress in self-improving LMMs, they still rely on structured intermediate signals such as reconstruction objectives, heuristic quality filtering, or semantic similarity-based scoring. Consequently, their self-evolution remains partially guided by external priors rather than emerging purely from the model’s own reasoning dynamics.
@@ 2 Related Work
In the context of large language models (LLMs), several recent works [5, 31, 6] aim to self-improve the model capabilities (e.g., scientific reasoning) without significant human annotations. The work of [5] proposes an approach to improve the implicit reasoning capabilities of the base model through self-play and majority-vote reward signals. Sheikh Shafayat et al. [31] explores self-consistency as a training reward within an reinforcement learning (RL) framework, demonstrating performance on synthetic reasoning task. Recently, the work of [6] introduces a self-play based method where three roles (Proposer, Solver, and Judge) are instantiated from the same base LLM and trained via RL to construct questions, answers, and the evaluation inter-play between them. While achieving promising results, these approaches assume a certain form of verifiable environment in the form of a judge [6] or heuristic and discrete rewards (e.g., majority vote) [5, 31] which can likely struggle with model collapse.
In the context of large multimodal models (LMMs), Li et al. [18] introduces a self-rewarding technique to improve visual understanding by decomposing it into perception and language-reasoning via explicit perception reward by the LMM itself. However, the self-reward mechanism relies on the SFT cold-start dataset to learn the desired format. The work of Zhang et al. [47] leverages image-level and instance-level reconstructions in a two-stage RL loop to iteratively improve perceptual capabilities. However, this image-and instance-level reconstruction of meta-data relies on an external model (e.g., OmniGen2 [40] and Qwen-Image [39]). Additionally, a recent work Vision-Zero [36], trains LMMs through a multi-agent social deduction game (“Who is the Spy?”) played directly on images. The framework introduces adversarial roles (spy vs. civilians), iterative clue and voting rounds, and alternates between self-play and verifiable reinforcement learning (RLVR) depending on whether the identity of the spy is known. Although Vision-Zero does not require explicit human annotations, it relies on constructing image pairs with subtle, controlled differences produced through automated rendering, image editing, or regeneration pipelines relying on external generative systems (GPT-4o [27] and Gemini [7]).
@@ 3.1 Unsupervised Self-Evolving LMM
In this work, we design an unsupervised self-evolving LMM approach (EvoLMM) motivated by the recent self-question-based technique (SQLM) [5] in the LLM domain. The SQLM improves an LLM by generating its own questions and reinforcing answer agreement. However, SQLM relies on a discrete majority-vote reward, which we empirically show to be sub-optimal in the multimodal setting where the LMM is required to ground its questions based on the input image and perform complex multimodal reasoning (e.g., visual math reasoning). We observe that early-stage Solver outputs over images are highly variable, leading to frequent zero-reward updates and unstable optimization. This prevents the model from making gradual progress when using discrete majority-vote reward.
@@
Training Optimization. In our framework, the Proposer and Solver are trained jointly in an online reinforcement-learning loop. Since no ground-truth labels or external scoring functions are available, both policies rely entirely on internally generated reward signals. To ensure stable learning, we adopt a REINFORCE policy gradient [38] with two key components: (i) moving-average baselines to reduce variance, and (ii) token-level KL regularization to monitor deviation from the pretrained base LMM.
@@ 3.2 EvoLMM Training Insights
To present further insights into our continuous self-consistency reward-based training, we conduct an experiment to post-train Qwen2.5-VL-7B [3] on ChartQA [23] training images (without any question-answer pairs or metadata). As shown in Figure 5, our self-consistency reward yields a graded score proportional to the empirical agreement among Solver outputs. Even partial agreement (e.g., 2 out of 5 consistent samples) produces a meaningful positive reward. Furthermore, our continuous feedback reduces collapse and accelerates Solver improvement (see Figure 4). The continuous Solver reward changes smoothly and slightly dips at partial consensus, while the Proposer peaks at moderate difficulty (see step 5122 in Figure 5).
@@ 4 Experiments
Vision-Zero (CLEVR) [36] $84.24$ $68.43$ $23.96$ $43.86$ $80.35$ $82.64$ $88.50$ $51.44$
Qwen2.5-VL-7B (Baseline) [3] $84.00$ $68.46$ $23.91$ $43.78$ $80.44$ $82.61$ $88.30$ $51.11$
@@ 4.1 Dataset
Our training uses only raw images, with no question–answer pairs or metadata. We sample roughly 1,000 images each from six widely used benchmarks such as ChartQA [23], AI2D [15], InfographicVQA [24], PlotQA [26], ChartX [41], and Geometry3K [2] totaling $\sim$6k images. These datasets cover charts, plots, scientific diagrams, and geometric figures, providing diverse visual contexts for the Proposer to generate grounded mathematical questions. This diversity encourages the Solver to develop stable, structured reasoning as the system self-evolves using internal consistency alone.
@@ 4.3 Evaluation
We evaluate our self-evolving framework on a broad suite of mathematical and visually grounded reasoning benchmarks, including ChartQA [23], MathVista [20], MathVision [35], MathVerse [49], InfoGraphic-VQA [24], AI2D [15], ScienceQA [21], and MMMU [44]. These datasets test a range of capabilities from chart interpretation and symbolic math to diagram understanding—and all metrics follow their standard accuracy protocols. We use the official evaluation splits and apply identical inference settings across models, ensuring that performance differences arise solely from our self-evolving training rather than task-specific tuning. All evaluations are run on AMD MI250X GPUs using the lmms-eval framework [48], with HuggingFace Transformers v4.38 and bfloat16 precision for consistency with training.
@@
Qwen2.5-VL-7B (Base) [3] $84.00$ $68.46$ $23.91$ $43.78$ $80.44$ $82.61$ $88.30$ $51.11$
@@
InternVL3-8B-Instruct (Base) [52] $82.40$ $65.2$ $25.36$ $31.62$ $68.77$ $83.19$ $97.77$ $52.78$
@@
Gemma3-12B-It (Base) [33] $55.64$ $60.13$ $24.53$ $28.96$ $50.69$ $79.05$ $83.89$ $48.11$
@@
Llama-3.2-11B-Vision-Instruct (Base) [25] $29.24$ $46.59$ $23.47$ $37.23$ $56.69$ $46.44$ $56.87$ $47.93$
@@ 4.4 Main Results
In Table 1, we compare our proposed EvoLMM framework against two baselines: (1) the original pretrained Qwen2.5-VL-7B model [3] without any additional training, and (2) a direct multimodal adaptation [5] that uses a discrete majority-vote reward during training. Qwen2.5-VL-7B (Baseline). The baseline model performs well on visually grounded reasoning tasks that require direct reading or surface-level interpretation (e.g., ChartQA trend identification, InfoGraphic-VQA object recognition). However, the model’s performance drops on tasks requiring multi-step inference or symbolic reasoning (e.g., MathVista, MathVerse), indicating that pretrained multimodal alignment alone is insufficient for deeper reasoning. We also report results from recent Vision-Zero [36], whose self-play formulation bears partial similarity to ours through autonomous multi-agent learning. However, Vision-Zero leverages synthetic image pairs generated using external models such as GPT-4o [27] and Gemini [7], introducing implicit supervision that falls outside our unsupervised setting. Vision-SR1 [18] differs entirely from our constraints, as it relies on paired question–answer supervision across both perception and reasoning stages.
@@ 4.5 Comparison Across Training Settings
We evaluate the self-evolving framework under three fine-tuning strategies such as LoRA [10], QLoRA [9], and full-parameter fine-tuning—while keeping the Proposer–Solver pipeline and continuous reward formulation identical across settings (see Table 2).
@@ 4.6 Self Evolution Across Different Backbones
To evaluate the generality of our approach, we apply the same Proposer–Solver training setup to several large multimodal models, including Qwen3-VL-8B-Instruct [3], InternVL3-8B [52], Gemma-3-12B-it [33], and Llama-3.2-11B-Vision-Instruct [25]. As shown in Table 3, in all cases, the models are trained without question–answer supervision, using only the proposed continuous self-consistency reward. We observe consistent improvements of roughly $2-3\%$ across all backbones, confirming that our method is architecture-agnostic and does not rely on specific dataset alignments or handcrafted supervision. Models with stronger visual grounding (e.g., InternVL3-8B [52]) show faster convergence in early training, while those with stronger linguistic reasoning (e.g., Gemma-3-12B [33]) display larger late-stage gains as solver reasoning stabilizes. These results demonstrate that continuous self-reward effectively generalizes across diverse LMM families. This shows that our reward mechanism enhances structured visual reasoning rather than surface-level pattern matching, providing a scalable and backbone-independent path toward autonomous multimodal reasoning.
@@ 5 Conclusion
Supplementary Material
This supplementary document provides expanded details, extended analyses, and deeper explanations supporting the main submission. We present (1) a complete training algorithm in  format, (2) comprehensive ablations with descriptive insights, (3) additional qualitative samples illustrating reasoning evolution, (4) curriculum diagnostics, (5) expanded backbone-scaling analysis, and (6) limitations and failure cases. All experimental results are derived from our EvoLMM: Self-Evolving Large Multimodal Models with Continuous Rewards
6 Overview
Our goal is to improve multimodal reasoning in LMMs without any human annotations, metadata, or external reward models. EvoLMM achieves this through a cooperative Proposer–Solver mechanism trained entirely using a continuous self-consistency reward. The training is performed over raw images only, making it fundamentally different from supervised or reward-distilled pipelines.
This supplementary provides additional clarity and extensions to strengthen the reproducibility and understanding of EvoLMM.
7 Detailed Training Algorithm
Below we provide the complete EvoLMM training loop in algorithmic format (see Algorithm 2). This version includes all variables, reward functions, and update rules described in the main paper. This formulation ensures stable co-adaptation between Proposer and Solver while preventing divergence from the pretrained backbone.
Output: Trained Proposer $\pi_{\phi}$ and Solver $\pi_{\theta}$
Initialize Proposer $\pi_{\phi}\leftarrow$ LoRA($\pi_{\text{ref}}$)
Initialize Solver $\pi_{\theta}\leftarrow$ LoRA($\pi_{\text{ref}}$)
Initialize EMA baselines $b_{\text{prop}},b_{\text{sol}}$
Set number of answer samples NN, entropy targets $(\mu_{H},\sigma_{H})$
for each training step do
  Sample image $x\sim\mathcal{X}$
 // Step 1: Proposer generates a question
  Generate question $q\sim\pi_{\phi}(q|x)$
 // Step 2: Solver answers the question multiple times
  Sample NN answers $\{y_{i}\}_{i=1}^{N}\sim\pi_{\theta}(\cdot|x,q)$
 // Step 3: Compute empirical answer distribution
  Estimate $p(a|x,q)$ and entropy $H(x,q)$
 // Step 4: Continuous Solver Reward
  for $i=1$ to NN do
     Compute agreement score $p(y_{i}|x,q)$
     Compute reward:
 // Step 5: Entropy-Based Proposer Reward
 // Step 6: Policy Gradient Updates
  Update Solver using REINFORCE:
 Update Proposer every $K$ steps:
$\nabla_{\phi}\sim(r^{\text{prop}}-b_{\text{prop}})\nabla\log\pi_{\phi}$
 Apply KL-regularization controller:
8 Expanded Ablation Experiments
8.1 Effect of Number of Solver Samples (N)
We evaluate $N\in\{3,5,8,12\}$. Below we expand the descriptive behavior of each setting:
N = 3 (Low-Compute, High Variance).
The reward signal becomes noisy due to insufficient samples to estimate a stable empirical answer distribution. This causes unstable early-stage training and leads the Proposer to generate semantically vague questions.
N = 5 (Optimal).
Our default choice. It provides a balanced, low-variance estimate of $p(a|x,q)$, enabling:
  • smoother reward gradients,
  • robust entropy estimation for the Proposer,
  • stable Solver convergence.
N = 8 (Marginal Gains).
Although reward curves become smoother, gains diminish while compute cost increases by $\sim 1.6\times$. The Solver also tends to become overconfident.
N = 12 (Over-Consensus Collapse).
Solver outputs collapse into near-deterministic samples, reducing reasoning diversity. The Proposer is forced toward overly difficult questions, breaking the entropy-based curriculum.
Conclusion: N=46N=4{-}6 is ideal. Excessively large $N$ harms training dynamics.
8.2 Effect of Backbone Size (3B → 72B)
We study Qwen2.5-VL backbones ranging from 3B to 72B. Larger models exhibit stronger absolute performance, but smaller models learn relatively more.
Small Models (3B/7B).
These models benefit greatly from self-evolution, often learning to structure their reasoning where none existed before.
Medium Models (32B).
Demonstrate more refined question understanding and stable compositional reasoning, benefiting from stronger visual grounding.
Large Models (72B).
Exhibit exceptional stability and produce highly consistent multi-step reasoning chains. Self-evolution enhances an already strong solver, resulting in high absolute improvements.
8.3 Ablation on Finetuning Strategies
LoRA (Best Overall).
LoRA preserves the backbone’s pretrained visual-linguistic alignment while enabling stable co-evolution. It produces the most reliable gains.
QLoRA.
Reduces memory footprint significantly. However, quantization noise slightly impairs reward sensitivity, leading to smaller improvements.
Full Fine-Tuning.
Highly unstable without external supervision. The Solver overfits quickly, and the Proposer often collapses into trivial repeated questions.
Key Insight: Parameter-efficient finetuning is essential for stable self-evolution.
9 Curriculum Emergence: Detailed Analysis
A central emergent property of EvoLMM is the self-generated curriculum guided by entropy-based Proposer rewards.
9.1 Phase 1: Early Exploration (Steps 0–1000)
Questions are often generic or weakly grounded. Solver responses are highly inconsistent. Yet the continuous reward ensures non-zero gradients allowing the system to progress.
9.2 Phase 2: Alignment Phase (Steps 1000–3000)
The Proposer starts grounding questions in actual image content. Solver responses become less random, shorter, and less verbose. Entropy begins concentrating around the mid-band, suggesting the system is entering productive learning.
9.3 Phase 3: Co-Evolution Phase (Steps 3000–6000)
The Proposer crafts increasingly intricate problems involving multi-step visual reasoning (e.g., path constraints, trend comparisons). The Solver shows stable reasoning chains and higher answer agreement.
This emergent progression occurs without any handcrafted difficulty scheduling.
10 Expanded Qualitative Results
We include textual descriptions of qualitative samples across training phases.
10.1 Early-Stage Examples
Example: Bar Chart (Step 200).
  • Proposer: “Count the shapes.”
  • Solver: {“2”, “D”, “Yes”, “C”, “2”}
The question is under-specified; Solver responses are unrelated. Reward remains low but non-zero, guiding early adaptation.
10.2 Mid-Stage Examples
Example: Line Plot (Step 2600).
  • Proposer: “Which year has the steepest decline?”
  • Solver: Mostly consistent answers (e.g., “2016”).
The Proposer demonstrates grounded numerical reasoning.
10.3 Late-Stage Examples
Example: Geometric Diagram (Step 5400).
  • Proposer: “From vertex P, which labeled vertex is reachable without crossing diagonal edges?”
  • Solver: {“R”, “R”, “R”, “S”, “R”}
Solver displays highly structured reasoning, aligning with underlying geometric constraints.
Model RefCOCOg DocVQA OCRBench Pope RealworldQA IFEval MME
Qwen-2.5-VL-7B 87.2 93.1 84.1 87.6 68.7 67.0 2318.8
+ EvoLMM (Ours) 88.0 94.4 85.0 88.4 69.9 69.0 2375.9
Table 5: Open-ended evaluation of baseline and our method, in a zero-shot setting, on additional multimodal benchmarks. Additional benchmark results demonstrating the effect of self-evolution on Qwen-2.5-VL-7B in zero-shot setting.
11 Scaling Behavior Across Backbones
Large models not only improve final accuracy but also:
  • generate higher-quality teaching-like questions,
  • maintain stable entropy distributions,
  • converge to deterministic reasoning,
  • exhibit deeper symbolic reasoning even in diagrammatic contexts.
This demonstrates that self-evolving training scales naturally without requiring reward retuning.
12 Failure Cases and Limitations
Ambiguous or Overly Abstract Questions.
The Proposer occasionally generates vague questions such as: “Which trend is overall the most stable?”
Perception Bottlenecks.
Thin lines, tiny grids, or dense legends sometimes mislead the Solver.
Stylistic Overfitting.
On chart-heavy datasets, the Proposer may overfit question templates.
Deterministic Solver Collapse (Large N).
Excessively large sample counts condition the Solver to produce nearly identical responses, reducing diversity and harming learning.
13 Additional Implementation Details
13.1 Training Configuration
[…]
Model (Ours - Base) ChartQA ScienceQA MathVista MMMU
InternVL3-1B ($\Delta$) +2.9% +2.3% +3.1% +2.1%
Qwen3-VL-2B ($\Delta$) +3.3% +2.5% +2.9% +2.6%
Table 6: Absolute performance gains on small baseline models (Intern-VL3-1B and Qwen3-VL-2B) using our self-evolution.
Category Baseline Ours $\Delta$
(Qwen-2.5-VL-7B) (EvoLMM)
Chart comprehension 82.5 86.8 +4.3
Symbolic reasoning 68.2 71.4 +3.2
Diagrammatic geometry 73.6 76.1 +2.5
Table 7: Performance comparison across different categories.
14 Additional Quantitative Results
14.1 Generalization to Open-Ended Benchmarks
Table 5 evaluates the self-evolved Qwen-2.5-VL-7B model on additional multimodal benchmarks in a strict zero-shot setting. Importantly, EvoLMM training does not use images or annotations from these datasets. Despite this, the self-evolved model consistently improves over the baseline across referring expression grounding (RefCOCOg), document understanding (DocVQA), OCR (OCRBench), real-world QA (RealworldQA), instruction following (IFEval), and overall multimodal evaluation (MME). These results indicate that the benefits of self-evolution extend beyond structured math reasoning tasks and generalize to more open-ended and deployment-oriented multimodal scenarios.
14.2 Small model behaviour
Table 6 reports absolute performance gains achieved through EvoLMM self-evolution on compact multimodal backbones (InternVL3-1B and Qwen3-VL-2B). Despite their limited capacity, both models consistently benefit from self-evolution across structured reasoning benchmarks, with improvements ranging from +2.1% to +3.3%. These results demonstrate that EvoLMM is not restricted to large-scale models, but remains effective for smaller architectures, highlighting its scalability and robustness across model sizes.
14.3 Fine-Grained Benchmark Analysis
Table. 7 report per-category scores on ChartQA, MathVista, and InfoGraphic-VQA. Our EvoLMM improves both data-reading and symbolic reasoning accuracy, with the largest gain (+4.1%) in visual numerical comparison tasks. We observe smaller yet consistent improvements in OCR-heavy subsets, confirming that self-consistency reward promotes structured reasoning rather than superficial text recognition.
15 Conclusion
This supplementary material expands the technical depth of EvoLMM and highlights how continuous reward design, entropy-driven curriculum emergence, and stable role decomposition enable LMMs to self-improve without external supervision. The detailed ablations and qualitative analyses reaffirm the robustness and generality of the proposed framework across model sizes, datasets, and training configurations.